A piece of iron of unknown mass has an initial temperature 210∘C. It is dropped into an aluminum container of mass 0.2 kg containing 1 litre (1.0 kg) of water both of which are at a temperature of temperature 20 ∘C . The final equilibrium temperature of the system when energy transfer between the iron and the water finally stops is 28.5 ∘C. (Assume no thermal energy gets lost.) What is the mass of the iron piece? Express your answer in kg to two decimal places.

The specific heat of iron is 450 J/kg ∘C, the specific heat of aluminum is 920 J/kg ∘C and the specific heat of water is 4186 J/kg ∘C.
Answer #1

heat lost by iron = heat gained by aluminum + water

or, misi(210C - 28.5C) = mwsw(28.5C - 20C) + masa(28.5C - 20C)

or, mi(450 J/kg ∘C)(210C - 28.5C) = (1 kg)(4186 J/kg ∘C)(28.5C - 20C) + (0.2 kg)(920 J/kg ∘C)(28.5C - 20C)

or, mi = 0.45 kg - mass of the iron piece

***********************************************************************************************
Check the answer and let me know immediately if you find something wrong or missing... I will rectify the mistakes asap if any